. Julia and John prefer more money to less and have transitive prefernces. Each of them faces the following decision problem. S260 $80 M T SX 30% 50% 20% S320 $80 $0 $320 S0 (a) Suppose that Julia is risk neutral. If she chooses A and then U, what can we deduce about the possible values of X and p? (b) Continue to assume that Julia is risk neutral. Suppose that X-250, p 0.9. Julia is given a choice between facing the above decision tree and playing the lottery $320 S260 $80 S60 What will she choose? If she chooses the decision tree 0.6 0.2 0. 0.1 explain what choices she makes, otherwise explain why she prefers the lottery (c) Suppose that X-100 and p 0.9 and John starts by choosing A. Then he says that he is indifferent between U and T. Is he risk neutral, risk averse or risk loving?
Answer #1

(a)The path AU gives Julia a wealth of $260 with certainty. Given that she is risk neutral, we can conclude the following:

(i) She gets less than $260 with certainty from path D

X < 260

(ii) If she follows path AT, she gets a wealth equal to or less than $260 (with uncertainty).

Using this information, we can calculate the value of p as follows:

320p+ 0(1-p) < 260

320p< 260

260 320

p<0.8125

(b) Given :

x -, p 0.9 250

From the decision tree, we have the following wealths from the different paths calculates as below:

Path D:

X-250

Path AU gives $260

Path AB gives 80

Path AM gives:

0.3 * 320 + 0.2 * 80 + 0.5 * 0 112

Path AT gives:

p*320 + (1-p)*0

= 0.9*320 + 0.1*0

288

From the lottery, Julia gets the following wealth:

0.6*320 + 0.2*260 + 0.1*80 + 0.1*60

=258

Given the calculations, we can conclude that Julia, bein a risk neutral wealth maximizer (prefers more money to less) would choose path AT from the decision tree because that maximizes her expected wealth with $288.

(c) Given:

X-100, p 0.9

Also, John is indifferent between U and T after he chooses A. So we have:

Sure wealth from U = $260

Expected wealth from T

=p*320 + (1-p)*0

=0.9*320 + 0.1*0

=288

John is a wealth maximizer who prefers more money to less. Also, he is risk neutral because he is indifferent between the outcomes.